Write the equation of the line that passes through point (-2,3)with a slope of -4.
Please help me

Answer by Edwin McCravy(20055)

Write the equation of the line that passes through point
(-2,3) with a slope of -4. 
Please help me

Rule:
--------------------------------------------------------------------

To write the equation of the line that passes through 
point (x1, y1) with a slope of m is gotten by plugging 
in the formula below and simplifying:

y - y1 = m(x - x1)

-------------------------------------------------------------------

In your case (x1, y1) = (-2, 3),
and m = -4, So x1 = -2 and y1 = 3

y - y1 = m(x - x1) becomes

 y - 3 = -4(x - (-2) )
 y - 3 = -4(x + 2)
 y - 3 = -4x - 8
     y = -4x - 5

Edwin